Y= tan^-1(x+1/2x+3)


1
Expert's answer
2022-07-03T12:44:25-0400
"y'=(\tan^{-1}(\dfrac{x+1}{2x+3}))'"

"=\dfrac{1}{1+(\dfrac{x+1}{2x+3})^2}(\dfrac{x+1}{2x+3})'"

"=\dfrac{(2x+3)^2}{4x^2+12x+9+x^2+2x+1}(\dfrac{2x+3-2(x+1)}{(2x+3)^2})"

"=\dfrac{1}{5x^2+14x+10}"